A block of weight 45.3 N is hanging from a rope. The tension from the rope is pulling upward on the block. The block is accelerating downward at a rate of 1.46 m/s2. What is the magnitude of the tension in the rope?
Answer #1

Let the tension be T

Accelerating downwards mean weight has a larger magnitude

So, W - T = m*a = W*a /g

T = W - W*a/g = 45.3 - [45.3*1.46/9.8] = 38.55 N

So tension in the rope is 38.55 N upwards
Answer #2

mg-T=ma => =45.3-T =(45.3/9.81)*1.46 =6.7419N taking g=9.81

=> T=45.3-6.7419 = 38.5581N